# The Rational Torsion Subgroup of $J_{0}(\mathfrak{p}^{r})$

### v1

@@ Abstract
Let $\mathfrak{n}=\mathfrak{p}^{r}$ be a prime-power ideal of $\mathbb{F}_{q}[T]$ with $r\geq 2$. We study the rational torsion subgroup $\mathcal{T}(\mathfrak{p}^{r})$ of the Drinfeld modular Jacobian $J_{0}(\mathfrak{p}^{r})$. We prove that the prime-to-$q(q-1)$ part of $\mathcal{T}(\mathfrak{p}^{r})$ is equal to that of the rational cuspidal divisor class group $\mathcal{C}(\mathfrak{p}^{r})$ of the Drinfeld modular curve $X_{0}(\mathfrak{p}^{r})$. As we completely computed the structure of $\mathcal{C}(\mathfrak{p}^{r})$, it also determines the structure of the prime-to-$q(q-1)$ part of $\mathcal{T}(\mathfrak{p}^{r})$.
@@ 1 Introduction
For a positive integer $N$, let $J_{0}(N)$ be the Jacobian variety of the classical modular curve $X_{0}(N)$ and $\mathcal{T}(N):=J_{0}(N)(\mathbb{Q})_{\text{tors}}$ its rational torsion subgroup. By the Mordell-Weil theorem, $\mathcal{T}(N)$ is a finite abelian group. Let $\mathcal{C}(N)$ be the rational cuspidal divisor class group of $X_{0}(N)$ defined in [16, p. 280]. By a theorem of Manin and Drinfeld, we have
@@
In the early 1970s, for any prime $p$, Ogg [10, Conj. 2] conjectured that $\mathcal{C}(p)=\mathcal{T}(p)$ and computed that $\mathcal{C}(p)$ is a cyclic group of order $\frac{p-1}{(p-1,12)}$. Later in 1977, Mazur [9, Thm. (1)] proved this conjecture by studying the Eisenstein ideal of the Hecke algebra of level $p$. There is a generalized Ogg’s conjecture:
@@
In [16, Thm. 1.7], Yoo et al. completely determined the structure of $\mathcal{C}(N)$ for arbitrary positive $N$. Consequently, if the above conjecture holds true, then the structure of $\mathcal{T}(N)$ is determined as well. The most notable current achievement in this regard is as follows:
Theorem 1.2 (Yoo [17, Thm. 1.4], Mazur, Lorenzini, Ling, Ohta, Ren, et al.).
@@
In [17, (1.1)], in a particular case, when $N=p^{r}$ is a prime power with $r\geq 2$, it has been established that $\mathcal{C}(p^{r})_{\ell}=\mathcal{T}(p^{r})_{\ell}$ holds for any prime $\ell\nmid 2p$. The condition on $\ell$ here is less restrictive compared to the general case.
@@
Let $\Gamma_{0}(\mathfrak{n})$ act on $\Omega$ via linear fractional transformations. Drinfeld proved in [1] that the quotient $\Gamma_{0}(\mathfrak{n})\backslash\Omega$ is the space of $\mathbb{C}_{\infty}$-points of an affine curve $Y_{0}(\mathfrak{n})$ defined over $K$, representing a moduli space of rank-$2$ Drinfeld modules. The unique smooth projective curve over $K$ containing $Y_{0}(\mathfrak{n})$ as an open subvariety is denoted by $X_{0}(\mathfrak{n})$, termed as the Drinfeld modular curve of level $\mathfrak{n}$. Let $J_{0}(\mathfrak{n})$ denote the Jacobian variety of $X_{0}(\mathfrak{n})$ and $\mathcal{T}(\mathfrak{n}):=J_{0}(\mathfrak{n})(K)_{\text{tors}}$ its rational torsion subgroup. According to the Lang-Néron theorem, $\mathcal{T}(\mathfrak{n})$ is a finite abelian group. To investigate the structure of $\mathcal{T}(\mathfrak{n})$, we study the rational cuspidal divisor class group $\mathcal{C}(\mathfrak{n})$ of $X_{0}(\mathfrak{n})$, as defined in [8]. By Gekeler [6, Thm. 1.2], $\mathcal{C}(\mathfrak{n})$ is a finite group, leading to the inclusion
@@
The structure of $\mathcal{C}(\mathfrak{n})$ is known in some special cases. First, when $\mathfrak{n}=\mathfrak{p}\lhd A$ is a prime ideal, the structure of $\mathcal{C}(\mathfrak{p})$ was fully determined by Gekeler [5, Cor. 3.23]. Second, Papikian and Wei [14, Thm. 1.3 and Prop. 1.5] determined the prime-to-$(q-1)$ part of $\mathcal{C}(\mathfrak{n})$ for any square-free ideal $\mathfrak{n}\lhd A$. In particular, when $\mathfrak{n}=\mathfrak{p}\mathfrak{q}$ is square-free with two prime ideals $\mathfrak{p}$ and $\mathfrak{q}$ of $A$, the structure of $\mathcal{C}(\mathfrak{p}\mathfrak{q})$ was entirely determined in [12, Thm. 1.1]. Third, when $\mathfrak{n}=\mathfrak{p}^{r}\lhd A$ is a prime-power ideal with $r\geq 2$, the structure of $\mathcal{C}(\mathfrak{p}^{r})$ was completely determined in the following:
Theorem 1.4 ([8, Thm. 3.5]).
@@
Moreover, an explicit basis of $\mathcal{C}(\mathfrak{p}^{r})$ is provided in [8, Thm. 3.5].
@@
When $\mathfrak{n}\lhd A$ is square-free or $\deg(\mathfrak{n})=3$, all the cusps of $X_{0}(\mathfrak{n})$ are rational by [13, Lem. 3.1]. Thus, the cuspidal divisor groups “$\mathcal{C}(\mathfrak{n})$” of $J_{0}(\mathfrak{n})$ defined in [14, p. 522] and [13, p. 399] in the above cases are the same as the rational cuspidal divisor class groups of $X_{0}(\mathfrak{n})$ defined in [8]. However, this phenomenon does not universally hold. For instance, when $\mathfrak{n}=(T^{2}+T+1)^{2}$ and $q=2$, the cuspidal divisor group of $J_{0}(\mathfrak{n})$ is strictly larger than the rational cuspidal divisor class group of $X_{0}(\mathfrak{n})$ by [13, Ex. 8.8] and Thm. 1.4.
The following presents an analogue of generalized Ogg’s conjecture:
Conjecture 1.6.
@@
This conjecture remains open, and the followings are some partial results:
@@
    Pál [11, Thm. 1.4] proved that $\mathcal{C}(\mathfrak{p})=\mathcal{T}(\mathfrak{p})$ for any prime ideal $\mathfrak{p}\lhd A$.
@@
    Papikian and Wei [13, Thm. 1.7] showed that $\mathcal{C}(\mathfrak{n})=\mathcal{T}(\mathfrak{n})$ for an ideal $\mathfrak{n}\lhd A$ generated by $T^{3}$ or $T^{2}(T-1)$.
@@
    Papikian and Wei [14, Thm. 1.3] demonstrated that if $\mathfrak{n}\lhd A$ is a square-free ideal, then $\mathcal{C}(\mathfrak{n})_{\ell}=\mathcal{T}(\mathfrak{n})_{\ell}$ for any prime $\ell$ not dividing $q(q-1)$.
@@
Let $\mathfrak{n}=\mathfrak{p}^{r}\lhd A$ be a prime-power ideal with $r\geq 2$. For any prime $\ell\nmid q(q-1)$, we have
@@
Both the square-free case discussed in [14, Thm. 1.3] and the prime-power case in Main Theorem utilize similar approaches in their proofs. However, a notable difference arises in the choice of tools: the former employs the Atkin-Lehner involutions $W_{\mathfrak{m}}$ for studying Eisenstein ideals, while the latter relies on the Hecke operator $U_{\mathfrak{p}}$ for the same purpose.
We provide a brief overview of the proof for Main Theorem. Fix a prime-power ideal $\mathfrak{n}=\mathfrak{p}^{r}\lhd A$ with $r\geq 2$. The Hecke algebra $\mathbb{T}(\mathfrak{n})$ of level $\mathfrak{n}$ is defined as the $\mathbb{Z}$-algebra generated by all Hecke operators $T_{\mathfrak{q}}$, where $\mathfrak{q}\lhd A$ are primes, acting on the group $\mathcal{H}_{0}(\mathfrak{n},\mathbb{Z})$ of $\mathbb{Z}$-valued $\Gamma_{0}(\mathfrak{n})$-invariant cuspidal harmonic cochains on the Bruhat-Tits tree $\mathscr{T}$ of $\operatorname{PGL}(2,K_{\infty})$. For further details, please refer to [13, p. 385], Sec. 3, and Sec. 4. The Eisenstein ideal $\mathfrak{E}(\mathfrak{n})$ of $\mathbb{T}(\mathfrak{n})$ is the ideal generated by the elements
@@
Additionally, $\mathbb{T}(\mathfrak{n})$ naturally acts on the Jacobian $J_{0}(\mathfrak{n})$ of $X_{0}(\mathfrak{n})$; see Sec. 5 and [13, p. 386] for more information. For the prime divisor $\mathfrak{p}$ of $\mathfrak{n}$, we usually write $U_{\mathfrak{p}}:=T_{\mathfrak{p}}\in\mathbb{T}(\mathfrak{n})$. Consider the following exact sequences of groups:
@@
Assuming the aboves and employing the short five lemma, we obtain Main Theorem. The first statement is achieved through explicit computation of the $U_{\mathfrak{p}}$ action on the cusps of $X_{0}(\mathfrak{p}^{r})$ combined with induction on $r$. To be specific, we decompose $U_{\mathfrak{p}}$ into $U_{\mathfrak{p}}=\alpha^{\ast}\circ\beta_{\ast}$, where $\alpha^{\ast}:J_{0}(\mathfrak{n}/\mathfrak{p})\rightarrow J_{0}(\mathfrak{n})$ and $\beta_{\ast}:J_{0}(\mathfrak{n})\rightarrow J_{0}(\mathfrak{n}/\mathfrak{p})$ are defined in Sec. 5. In Lem. 5.3, we demonstrate that $\beta_{\ast}(\mathcal{C}(\mathfrak{p}^{r}))_{\ell}=\mathcal{C}(\mathfrak{p}^{r-1})_{\ell}$ for any prime $\ell\neq p$. Assume the induction hypothesis that $\mathcal{C}(\mathfrak{p}^{r-1})_{\ell}=\mathcal{T}(\mathfrak{p}^{r-1})_{\ell}$ for any prime $\ell\nmid q(q-1)$. Note that $\mathcal{C}(\mathfrak{p})=\mathcal{T}(\mathfrak{p})$ serves as the initial condition for the induction. Then for any prime $\ell\nmid q(q-1)$,
@@
The second statement is established through an Atkin-Lehner type result; see Thm. 6.5. To be precise, for a coefficient ring $R$ (see Def. 3.2), let $\mathcal{E}(\mathfrak{n},R)[U_{\mathfrak{p}}]$ represent the group of $R$-valued $\Gamma_{0}(\mathfrak{n})$-invariant harmonic cochains on $\mathscr{T}$ that are annihilated by both the actions of the Eisenstein ideal $\mathfrak{E}(\mathfrak{n})$ and the Hecke operator $U_{\mathfrak{p}}$. We construct an explicit element $E_{\mathfrak{n}}$ in $\mathcal{E}(\mathfrak{n},R)[U_{\mathfrak{p}}]$; cf. Eq. (7) and Lem. 6.3. Utilizing Thm. 6.5, we demonstrate that $E_{\mathfrak{n}}$ generates $\mathcal{E}(\mathfrak{n},R)[U_{\mathfrak{p}}]$ over $R$. Considering the values of $E_{\mathfrak{n}}$ on specific edges of $\mathscr{T}$, we observe that the subgroup $\mathcal{E}_{0}(\mathfrak{n},R)[U_{\mathfrak{p}}]$ of $\mathcal{E}(\mathfrak{n},R)[U_{\mathfrak{p}}]$, consisting of cuspidal harmonic cochains, is contained in $R\left[M(\mathfrak{p})\right]\cdot E_{\mathfrak{n}}$, where $R\left[M(\mathfrak{p})\right]:=\{c\in R\mid M(\mathfrak{p})\cdot c=0\}$. By [12, Lem. 7.2], let $R:=\mathbb{Z}/\ell^{k}\mathbb{Z}$ for a prime $\ell\nmid q(q-1)$ and a sufficiently large $k$, then there exists an injective map:
$\mathcal{C}(\mathfrak{p}^{r})_{\ell}[U_{\mathfrak{p}}]\subseteq\mathcal{T}(\mathfrak{p}^{r})_{\ell}[U_{\mathfrak{p}}]\hookrightarrow\mathcal{E}_{0}(\mathfrak{p}^{r},R)[U_{\mathfrak{p}}]\subseteq R\left[M(\mathfrak{p})\right]\cdot E_{\mathfrak{n}}\cong\frac{\mathbb{Z}_{\ell}}{M(\mathfrak{p})\mathbb{Z}_{\ell}}.$ (1)
Moreover, according to Thm. 1.4 and Sec. 5, we construct explicitly a cyclic subgroup of $\mathcal{C}(\mathfrak{n})_{\ell}[U_{\mathfrak{p}}]$ that is isomorphic to $\mathbb{Z}_{\ell}/M(\mathfrak{p})\mathbb{Z}_{\ell}$; see Eq. (4), Sequence (6), and Prop. 6.9. Hence, we demonstrate that $\mathcal{C}(\mathfrak{p}^{r})_{\ell}[U_{\mathfrak{p}}]=\mathcal{T}(\mathfrak{p}^{r})_{\ell}[U_{\mathfrak{p}}]$.
@@
The idea of considering Eq. (3) and Sequence (1) comes from the proof of $\mathcal{C}(\mathfrak{p})=\mathcal{T}(\mathfrak{p})$ for the prime level case in [11, Thm. 1.4]. More precisely, Pál showed that the image of the $\ell$-primary part of $\mathcal{T}(\mathfrak{p})$, for a prime $\ell\neq p$, in the group $\Phi_{\infty}$ of connected components of the Néron model of $J_{0}(\mathfrak{p})$ at $\infty$ with respect to specialization, injects into $\mathcal{E}_{0}(\mathfrak{p},\mathbb{Z}/\ell^{k}\mathbb{Z})$ for large enough $k$; cf. [11, p. 133], which leads to [11, Thm. 7.19].
@@
The primes $\ell$ dividing $q(q-1)$ lie beyond the scope of our current approach. Considering a non-zero ideal $\mathfrak{n}\lhd A$, $J_{0}(\mathfrak{n})$ has bad reduction both at $\infty$ and at the primes dividing $\mathfrak{n}$. In particular, $J_{0}(\mathfrak{n})$ has split toric reduction at $\infty$. In instances where $\ell$ divides $q(q-1)$, the specialization map $\mathcal{T}(\mathfrak{n})_{\ell}\rightarrow(\Phi_{\infty})_{\ell}$ might not ensure injectivity; cf. [12, Lem. 7.2].
The paper is outlined in the following. We begin in Sec. 2 with an introduction to the rational cuspidal divisor class group of $X_{0}(\mathfrak{n})$. Sec. 3 covers harmonic cochains on the Bruhat-Tits tree $\mathscr{T}$ of $\operatorname{PGL}(2,K_{\infty})$ and their Fourier expansions. Following this, Sec. 4 discusses Hecke operators acting on the space of $R$-valued harmonic cochains on $\mathscr{T}$, where $R$ is a commutative ring with unity, alongside Eisenstein ideals for prime-power levels. Sec. 5 explores the action of Hecke operators on the Jacobian $J_{0}(\mathfrak{n})$ of the Drifeld modular curve $X_{0}(\mathfrak{n})$. Finally, all introduced tools are integrated into the proof of Main Theorem.

### v2

@@ Abstract
Let $\mathfrak{n}=\mathfrak{p}^{r}$ be a prime power ideal of $\mathbb{F}_{q}[T]$ with $r\geq 2$. We study the rational torsion subgroup $\mathcal{T}(\mathfrak{p}^{r})$ of the Drinfeld modular Jacobian $J_{0}(\mathfrak{p}^{r})$. We prove that the prime-to-$q(q-1)$ part of $\mathcal{T}(\mathfrak{p}^{r})$ is equal to that of the rational cuspidal divisor class group $\mathcal{C}(\mathfrak{p}^{r})$ of the Drinfeld modular curve $X_{0}(\mathfrak{p}^{r})$. As we completely computed the structure of $\mathcal{C}(\mathfrak{p}^{r})$, it also determines the structure of the prime-to-$q(q-1)$ part of $\mathcal{T}(\mathfrak{p}^{r})$.
Department of Mathematics, The Pennsylvania State University, State College, PA 16801, USA
E-mail address: kevinho@psu.edu
@@ 1 Introduction
For a positive integer $N$, let $J_{0}(N)$ be the Jacobian variety of the classical modular curve $X_{0}(N)$ and $\mathcal{T}(N):=J_{0}(N)(\mathbb{Q})_{\text{tors}}$ its rational torsion subgroup. By the Mordell-Weil theorem, $\mathcal{T}(N)$ is a finite abelian group. Let $\mathcal{C}(N)$ be the rational cuspidal divisor class group of $X_{0}(N)$ defined in [17, p. 280]. By a theorem of Manin and Drinfeld, we have
@@
In the early 1970s, for any prime $p$, Ogg [11, Conjecture 2] conjectured that $\mathcal{C}(p)=\mathcal{T}(p)$ and computed that $\mathcal{C}(p)$ is a cyclic group of order $\frac{p-1}{(p-1,12)}$. Later in 1977, Mazur [10, Theorem (1)] proved this conjecture by studying the Eisenstein ideal of the Hecke algebra of level $p$. There is a generalized Ogg’s conjecture:
@@
In [17, Theorem 1.7], Yoo et al. completely determined the structure of $\mathcal{C}(N)$ for arbitrary positive $N$. Consequently, if the above conjecture holds true, then the structure of $\mathcal{T}(N)$ is determined as well. The most notable current achievement in this regard is as follows:
Theorem 1.2 (Yoo [18, Theorem 1.4], Mazur, Lorenzini, Ling, Ohta, Ren, et al.).
@@
In [18, (1.1)], in a particular case, when $N=p^{r}$ is a prime power with $r\geq 2$, it has been established that $\mathcal{C}(p^{r})_{\ell}=\mathcal{T}(p^{r})_{\ell}$ holds for any prime $\ell\nmid 2p$. The condition on $\ell$ here is less restrictive compared to the general case.
@@
Let $\Gamma_{0}(\mathfrak{n})$ act on $\Omega$ via linear fractional transformations. Drinfeld proved in [1] that the quotient $\Gamma_{0}(\mathfrak{n})\backslash\Omega$ is the space of $\mathbb{C}_{\infty}$-points of an affine curve $Y_{0}(\mathfrak{n})$ defined over $K$, representing a moduli space of rank-$2$ Drinfeld modules. The unique smooth projective curve over $K$ containing $Y_{0}(\mathfrak{n})$ as an open subvariety is denoted by $X_{0}(\mathfrak{n})$, termed as the Drinfeld modular curve of level $\mathfrak{n}$. Let $J_{0}(\mathfrak{n})$ denote the Jacobian variety of $X_{0}(\mathfrak{n})$ and $\mathcal{T}(\mathfrak{n}):=J_{0}(\mathfrak{n})(K)_{\text{tors}}$ its rational torsion subgroup. According to the Lang-Néron theorem, $\mathcal{T}(\mathfrak{n})$ is a finite abelian group. To investigate the structure of $\mathcal{T}(\mathfrak{n})$, we study the rational cuspidal divisor class group $\mathcal{C}(\mathfrak{n})$ of $X_{0}(\mathfrak{n})$, as defined in [9]. By Gekeler [6, Theorem 1.2], $\mathcal{C}(\mathfrak{n})$ is a finite group, leading to the inclusion
@@
The structure of $\mathcal{C}(\mathfrak{n})$ is known in some special cases. First, when $\mathfrak{n}=\mathfrak{p}\lhd A$ is a prime ideal, the structure of $\mathcal{C}(\mathfrak{p})$ was fully determined by Gekeler [5, Corollary 3.23]. Second, Papikian and Wei [15, Theorem 1.3 and Proposition 1.5] determined the prime-to-$(q-1)$ part of $\mathcal{C}(\mathfrak{n})$ for any square-free ideal $\mathfrak{n}\lhd A$. In particular, when $\mathfrak{n}=\mathfrak{p}\mathfrak{q}$ is square-free with two prime ideals $\mathfrak{p}$ and $\mathfrak{q}$ of $A$, the structure of $\mathcal{C}(\mathfrak{p}\mathfrak{q})$ was entirely determined in [13, Theorem 1.1]. Third, when $\mathfrak{n}=\mathfrak{p}^{r}\lhd A$ is a prime power ideal with $r\geq 2$, the structure of $\mathcal{C}(\mathfrak{p}^{r})$ was completely determined in the following:
Theorem 1.4 ([9, Theorem 3.5]).
@@
Moreover, an explicit basis of $\mathcal{C}(\mathfrak{p}^{r})$ is provided in [9, Theorem 3.5].
@@
When $\mathfrak{n}\lhd A$ is square-free or $\deg(\mathfrak{n})=3$, all the cusps of $X_{0}(\mathfrak{n})$ are rational by [7, Proposition 6.7] and [14, Lemma 3.1]. Thus, the cuspidal divisor groups of $J_{0}(\mathfrak{n})$ defined in [15, p. 522] and [14, p. 399] in the above cases are the same as the rational cuspidal divisor class groups of $X_{0}(\mathfrak{n})$ defined in [9]. However, this phenomenon does not universally hold. For instance, when $\mathfrak{n}=(T^{2}+T+1)^{2}$ and $q=2$, the cuspidal divisor group of $J_{0}(\mathfrak{n})$ is strictly larger than the rational cuspidal divisor class group of $X_{0}(\mathfrak{n})$ by [14, Example 8.8] and Theorem 1.4.
There is an analogue of the generalized Ogg’s conjecture:
Conjecture 1.6 ([9, Conjecture 1.2]).
@@
This conjecture remains open, and the following are some partial results:
@@
    Pál [12, Theorem 1.4] proved that $\mathcal{C}(\mathfrak{p})=\mathcal{T}(\mathfrak{p})$ for any prime ideal $\mathfrak{p}\lhd A$.
@@
    Papikian and Wei [14, Theorem 1.7] demonstrated that $\mathcal{C}(\mathfrak{n})=\mathcal{T}(\mathfrak{n})$ for an ideal $\mathfrak{n}\lhd A$ generated by $T^{3}$ or $T^{2}(T-1)$.
@@
    Papikian and Wei [15, Theorem 1.3] showed that if $\mathfrak{n}\lhd A$ is square-free, then $\mathcal{C}(\mathfrak{n})_{\ell}=\mathcal{T}(\mathfrak{n})_{\ell}$ for any prime $\ell$ not dividing $q(q-1)$.
@@
Let $\mathfrak{n}=\mathfrak{p}^{r}\lhd A$ be a prime power ideal with $r\geq 2$. For any prime $\ell\nmid q(q-1)$, we have
@@
Both the square-free case discussed in [15, Theorem 1.3] and the prime power case in the Main Theorem utilize similar approaches in their proofs. However, a notable difference arises in the choice of tools: the former employs the Atkin-Lehner involutions $W_{\mathfrak{m}}$ for studying Eisenstein ideals, while the latter relies on the Hecke operator $U_{\mathfrak{p}}$ for the same purpose.
We provide a brief overview of the proof for the Main Theorem. Fix a prime power ideal $\mathfrak{n}=\mathfrak{p}^{r}\lhd A$ with $r\geq 2$. The Hecke algebra $\mathbb{T}(\mathfrak{n})$ of level $\mathfrak{n}$ is defined as the $\mathbb{Z}$-algebra generated by all Hecke operators $T_{\mathfrak{q}}$, where $\mathfrak{q}\lhd A$ are primes (including $\mathfrak{q}=\mathfrak{p}$), acting on the group $\mathcal{H}_{0}(\mathfrak{n},\mathbb{Z})$ of $\mathbb{Z}$-valued $\Gamma_{0}(\mathfrak{n})$-invariant cuspidal harmonic cochains on the Bruhat-Tits tree $\mathscr{T}$ of $\operatorname{PGL}(2,K_{\infty})$. For further details, please refer to [14, p. 385], Section 3, and Section 4. The Eisenstein ideal $\mathfrak{E}(\mathfrak{n})$ of $\mathbb{T}(\mathfrak{n})$ is the ideal generated by the elements
@@
Additionally, $\mathbb{T}(\mathfrak{n})$ naturally acts on the Jacobian $J_{0}(\mathfrak{n})$ of $X_{0}(\mathfrak{n})$; see Section 5 and [14, p. 386] for more information. For the prime divisor $\mathfrak{p}$ of $\mathfrak{n}$, we usually write $U_{\mathfrak{p}}:=T_{\mathfrak{p}}\in\mathbb{T}(\mathfrak{n})$. Consider the following exact sequences of groups:
@@
Assuming the above and employing the short five lemma, we obtain the Main Theorem. The first statement is achieved through explicit computation of the $U_{\mathfrak{p}}$ action on the cusps of $X_{0}(\mathfrak{p}^{r})$ combined with induction on $r$. To be specific, we decompose $U_{\mathfrak{p}}$ into $U_{\mathfrak{p}}=\alpha^{\ast}\circ\beta_{\ast}$, where $\alpha^{\ast}:J_{0}(\mathfrak{n}/\mathfrak{p})\rightarrow J_{0}(\mathfrak{n})$ and $\beta_{\ast}:J_{0}(\mathfrak{n})\rightarrow J_{0}(\mathfrak{n}/\mathfrak{p})$ are defined in Section 5. In Lemma 5.3, we demonstrate that $\beta_{\ast}(\mathcal{C}(\mathfrak{p}^{r}))_{\ell}=\mathcal{C}(\mathfrak{p}^{r-1})_{\ell}$ for any prime $\ell\neq p$. Assume the induction hypothesis that $\mathcal{C}(\mathfrak{p}^{r-1})_{\ell}=\mathcal{T}(\mathfrak{p}^{r-1})_{\ell}$ for any prime $\ell\nmid q(q-1)$. Note that $\mathcal{C}(\mathfrak{p})=\mathcal{T}(\mathfrak{p})$ serves as the initial condition for the induction. Then for any prime $\ell\nmid q(q-1)$,
@@
The second statement is established through an Atkin-Lehner type result; see Theorem 6.5. To be precise, for a coefficient ring $R$ (see Definition 3.2), let $\mathcal{E}(\mathfrak{n},R)[U_{\mathfrak{p}}]$ represent the group of $R$-valued $\Gamma_{0}(\mathfrak{n})$-invariant harmonic cochains on $\mathscr{T}$ that are annihilated by both the actions of the Eisenstein ideal $\mathfrak{E}(\mathfrak{n})$ and the Hecke operator $U_{\mathfrak{p}}$. We construct an explicit element $E_{\mathfrak{n}}$ in $\mathcal{E}(\mathfrak{n},R)[U_{\mathfrak{p}}]$; cf. Equation (7) and Lemma 6.3. Utilizing Theorem 6.5, we demonstrate that $E_{\mathfrak{n}}$ generates $\mathcal{E}(\mathfrak{n},R)[U_{\mathfrak{p}}]$ over $R$. Considering the values of $E_{\mathfrak{n}}$ on specific edges of $\mathscr{T}$, we observe that the subgroup $\mathcal{E}_{0}(\mathfrak{n},R)[U_{\mathfrak{p}}]$ of $\mathcal{E}(\mathfrak{n},R)[U_{\mathfrak{p}}]$, consisting of cuspidal harmonic cochains, is contained in $R\left[M(\mathfrak{p})\right]\cdot E_{\mathfrak{n}}$, where $R\left[M(\mathfrak{p})\right]:=\{c\in R\mid M(\mathfrak{p})\cdot c=0\}$. Let $R:=\mathbb{Z}/\ell^{k}\mathbb{Z}$ for a prime $\ell\nmid q(q-1)$ and a sufficiently large $k$. By [13, Lemma 7.2], there exists an injective map:
$\mathcal{C}(\mathfrak{p}^{r})_{\ell}[U_{\mathfrak{p}}]\subseteq\mathcal{T}(\mathfrak{p}^{r})_{\ell}[U_{\mathfrak{p}}]\hookrightarrow\mathcal{E}_{0}(\mathfrak{p}^{r},R)[U_{\mathfrak{p}}]\subseteq R\left[M(\mathfrak{p})\right]\cdot E_{\mathfrak{n}}\hookrightarrow\frac{\mathbb{Z}_{\ell}}{M(\mathfrak{p})\mathbb{Z}_{\ell}}.$ (1)
Moreover, according to Theorem 1.4 and Section 5, we construct explicitly a cyclic subgroup of $\mathcal{C}(\mathfrak{n})_{\ell}[U_{\mathfrak{p}}]$ that is isomorphic to $\mathbb{Z}_{\ell}/M(\mathfrak{p})\mathbb{Z}_{\ell}$; see Equation (4), Sequence (6), and Proposition 6.9. Hence, we demonstrate that $\mathcal{C}(\mathfrak{p}^{r})_{\ell}[U_{\mathfrak{p}}]=\mathcal{T}(\mathfrak{p}^{r})_{\ell}[U_{\mathfrak{p}}]$.
@@
The idea of considering Map (3) and Sequence (1) comes from the proof of $\mathcal{C}(\mathfrak{p})=\mathcal{T}(\mathfrak{p})$ for the prime level case in [12, Theorem 1.4]. More precisely, Pál showed that the image of the $\ell$-primary part of $\mathcal{T}(\mathfrak{p})$, for a prime $\ell\neq p$, in the group $\Phi_{\infty}$ of connected components of the Néron model of $J_{0}(\mathfrak{p})$ at $\infty$ with respect to specialization, injects into $\mathcal{E}_{0}(\mathfrak{p},\mathbb{Z}/\ell^{k}\mathbb{Z})$ for large enough $k$; cf. [12, p. 133], which leads to [12, Theorem 7.19].
@@
The primes $\ell$ dividing $q(q-1)$ lie beyond the scope of our current approach. Considering a non-zero ideal $\mathfrak{n}\lhd A$, $J_{0}(\mathfrak{n})$ has bad reduction both at $\infty$ and at the primes dividing $\mathfrak{n}$. In particular, $J_{0}(\mathfrak{n})$ has split toric reduction at $\infty$. In instances where $\ell$ divides $q(q-1)$, the specialization map $\mathcal{T}(\mathfrak{n})_{\ell}\rightarrow(\Phi_{\infty})_{\ell}$ might not ensure injectivity; cf. [13, Lemma 7.2].
The paper is outlined in the following. We begin in Section 2 with an introduction to the rational cuspidal divisor class group of $X_{0}(\mathfrak{n})$. Section 3 covers harmonic cochains on the Bruhat-Tits tree $\mathscr{T}$ of $\operatorname{PGL}(2,K_{\infty})$ and their Fourier expansions. Following this, Section 4 discusses Hecke operators acting on the space of $R$-valued harmonic cochains on $\mathscr{T}$, where $R$ is a commutative ring with unity, alongside Eisenstein ideals for prime power levels. Section 5 explores the action of Hecke operators on the Jacobian $J_{0}(\mathfrak{n})$ of the Drifeld modular curve $X_{0}(\mathfrak{n})$. Finally, all introduced tools are integrated into the proof of the Main Theorem.